# Path-based packing of icosahedral shells into multi-component aggregates

### v2

@@ Acknowledgments
    1. 1.
[…]
    2. 2.
      particles close to the vertices (marked in blue in the figure) rotate around the closest vertex counterclockwise of an angle $\theta_{CK}-30^{\circ}$; in this way, they reach the medians of the icosahedral facet;
    3. 3.
      the remaining particles (marked in yellow in the figure) rotate clockwise around the center of the closest facet edge, of an angle $60^{\circ}-\theta_{CK}$; in this way, they reach the edge of the facet.
    Ch2 shells with $k=2$, $k<h$ transform into the same AM2 shells; the transformation is of the same type, with opposite rotation directions, and rotation angles of $\theta_{CK}$ instead of $60^{\circ}-\theta_{CK}$, and $30^{\circ}-\theta_{CK}$ instead of $\theta_{CK}-30^{\circ}$.
We note that such transformations do not allow to obtain the optimal AM arrangement, since neighbouring particles rotating around different axes end up with inter-particle distances that are shorter than the ideal one. This occurs because the radii of the perfect Ch and AM shells with the same number of particles are different, as we show in the following.
The edge of the AM shell with unit inter-particle distance and AM indexes $p$, $q$ can be easily calculated by looking at Supplementary Figure 9. It corresponds to the side of the inner triangle, which is $p-1$, plus twice the projection of the segment between the vertices of the facet and of the inner triangle (which is $q+1$ long) on the facet edge:
$l_{AM}^{(p,q)}=p-1+2(q+1)\frac{\sqrt{3}}{2}=p+\sqrt{3}q+\sqrt{3}-1.$ (S77)
We write the edge length as a function of the CK indexes of the corresponding CK shell, by substituting $p=k-h+1$ and $q=h-1$:
$l_{AM}^{(h,k)}=(\sqrt{3}-1)h+k,$ (S78)
We recall that the edge length of the CK shell is
$l_{CK}^{(h,k)}=\sqrt{T^{(h,k)}}=\sqrt{h^{2}+k^{2}+hk}.$ (S79)
It is easy to verify that, for each couple of indexes $(h,k)$, $l_{AM}^{(h,k)}>l_{CK}^{(h,k)}$. The same holds for the radii, which are calculated by multiplying the edge length by $\sin{(2\pi/5)}$. CK shells are more compact than the corresponding AM ones, as particles are more densely packed: in CK shells, all particles are closed-packed (we recall that CK shells are obtained by folding an hexagonal closed-packed lattice), whereas in AM shells only a portion of particles is in closed-packed arrangement, the other ones forming only 5 or 4 nearest-neighbours bonds.
We can estimate the optimal size mismatch for icosahedra with AM shells by using the same geometric consideration made for CK icosahedra. Indeed, we can use the formula in Eq. S53 and S67, in which we shall replace $\sqrt{T}$ by the length of the AM edge in Eq. S78. As an example, we calculate the optimal size mismatch for adding a AM1 shell on a MC core, i.e. between the $(0,k)$ and $(1,k)^{\ast}$ icosahedral shells, with $k\geq 1$. The length of the edge of the $(1,k)^{\ast}$ shell is
$l_{AM}^{(1,k)}=k+\sqrt{3}-1,$ (S80)
and therefore the optimal size mismatch is given by
$sm_{MC-AM1}^{(k)}=\frac{2\,(1+\varepsilon_{MC})\sin{\left(\frac{2\pi}{5}\right)}k+1}{2\sin{\left(\frac{2\pi}{5}\right)}(k+\sqrt{3}-1)-1}-1.$ (S81)
Values are reported in Figure 7a of the main text.
S1.6 Comparison of the stability of AM1 and Ch1 shells
The data on the optimal mismatch in Lennard-Jones and Morse clusters have been obtained by calculating the energy, after local minimization, as a function of the mismatch for the different types of structures, and looking for the mismatch at which the energy is minimum. Two examples are reported in Supplementary Figure 10, where we consider MC, AM1 and Ch1 shells on Mackay cores made of 5 and 8 MC shells. The MC structures are better stabilised in a low-mismatch range. The AM1 shells find their optimal mismatch in an intermediate range and after a threshold mismatch they become unstable, i.e. they are not even local minima. On the contrary, the Ch1 structures are stable in a high-mismatch range and become unstable below a threshold mismatch. In the cases reported in Supplementary Figure 10 the range of mismatch values in which both AM1 and Ch1 structures are local minima is very narrow. In Supplementary Figure 10a we have reported the same data of Figure 3a of the main text for the addition of either a MC or a Ch1 shell on a five-shell Mackay core; data for the addition of a AM1 shell have been included, in order to compare the three possible arrangements of the outer shell.
Figure 10: Stability of MC, AM1 and Ch1 shells in Lennard-Jones clusters.
Binding energy per particle (in $\varepsilon$ units) as a function of the size mismatch of a an icosahedron of $i=5$ MC shells covered by single MC, AM1 and Ch1 shells and b an icosahedron of $i=8$ MC shells covered by single MC, AM1 and Ch1 shells.
Figure 11: Optimal mismatch for two AM shells on a MC core.
A core terminated by a $(1,k)^{*}$ AM1 shell is covered by either a $(1,k+1)^{*}$ or a $(2,k)^{*}$ shell. The optimal mismatch is evaluated according to Eq. S82 and Eq. S83.
Figure 12: Stability of two AM shells on a MC core.
a-b The two types of reconstruction undergone by the outer AM1 shell spontaneously upon local minimization, involving rotations of particle rings around all vertices. c Energy (in $\varepsilon$ units) as a function of the mismatch for Lennard-Jones clusters made of 5 Mackay shells upon which two shells are added: two Ch1 shells; two AM1 shells; an AM1 shell plus a shell reconstructed as in a; an AM1 shell plus a shell reconstructed as in b.
S1.7 Packing of AM shells
In the main text we have evaluated the optimal mismatch for adding a $(1,k)^{*}$ AM1 shell on a MC core terminated by a $(0,k)$ shell. Here we consider the possibility of adding a further shell of AM type. According to the path rules, the shell can be either of $(1,k+1)^{*}$ or $(2,k)^{*}$ type, i.e. AM1 or AM2. The optimal mismatch can be analytically evaluated as explained in the previous Sections, obtaining the formula below:
$sm_{AM1-AM1}^{(k)}=\frac{2\sin{\left(\frac{2\pi}{5}\right)}(k+\sqrt{3}-1)+1}{2\sin{\left(\frac{2\pi}{5}\right)}(k+\sqrt{3})-1}-1,$ (S82)
$sm_{AM1-AM2}^{(k)}=\frac{2\sin{\left(\frac{2\pi}{5}\right)}(k+\sqrt{3}-1)+1}{2\sin{\left(\frac{2\pi}{5}\right)}(k+2\sqrt{3})-2}-1.$ (S83)
Values of the optimal mismatch are reported in Supplementary Figure 11. As in the case of chiral shells (see Figure 3d in the main text), the optimal mismatch is larger for the class-changing step $(1,k)^{*}\to(2,k)^{*}$. This may lead to the conclusion that the outer shell should be preferentially of AM1 type (i.e. $(1,k+1)^{*}$) if the mismatch is zero or low. But this is not the case, as we show in Supplementary Figure 12. In fact the mismatch range in which the configuration with two AM1 shells is a local minimum is very narrow, and even when locally stable, its energy is quite high. Outside that range, the outer shell spontaneously reconstructs upon local minimization, as shown in Supplementary Figure 12a-b. This is due to the very unfavourable coordination of particles on the facet edges of the outer shell, which sit on only two nearest neighbours of the shell below (this point will be further discussed in the last Section). Both reconstructions improve that coordination by displacing atoms on fourfold sites of the substrate.
In Supplementary Figure 12c we calculate the energies of unreconstructed and reconstructed configurations of Lennard-Jones clusters. The reconstruction of Supplementary Figure 12b is more favourable, as it displaces a larger number of particles on sites of higher coordination with the substrate. Due to the instability of the configuration with two AM1 shells, the natural growth sequences lead to the formation of an outer AM2 shell, as shown in Figure 7b-c of the main text and discussed in the Supplementary Note 3.
Finally, we note that the configuration with two Ch1 shells, which has the same number of particles, is much more favourable in a wide range of mismatch, as shown in Supplementary Figure 12c. In fact, from the results in Supplementary Figure 12c it turns out also that the stability range of two Ch1 shells is enlarged with respect to that of one Ch1 shell (see Supplementary Figure 10a).
Appendix S2 Density Functional Theory (DFT) calculations
Figure 13: Clusters of the DFT calculations.
The structures considered in the calculations are of two types, with either $i=3$ or $i=4$ shells. In this figure we show only those containing BG and Ch1 shells, while the analogous structures containing 3 or 4 MC shells are not shown. The first structure ($i=3$) contains the central atom (0,0), the (0,1) MC shell, and the (1,1) BG shell, corresponding to the left and middle panels in a, for a total of 45 atoms. The second structure ($i=4$) contains in addition the (1,2) Ch1 shell, as shown in the right panel of a, for a total of 117 atoms. In b-e the different types of atom pair swaps are shown: in b and c between atoms of the (0,1) and (1,1) shells; in d and e between atoms of the (1,1) and (1,2) shells.
In Supplementary Tables 2-4 we report the data for the atomic pair exchanges shown in Supplementary Figure 13 for different types of alkali metal clusters. DFT [51] and Gupta potential [52] results are reported. The results have been obtained as explained in the Methods section of the main text. Here we recall that the energy differences between the configuration after and before the exchange are related to locally minimized configurations.
All structures contain a core made of $(0,0)$ and $(0,1)$ MC shells, upon which further shells are added, either of Ch1 class (the $(1,1)$ BG shell and the $(1,2)$ chiral shell) or of Ch0 class (the $(0,2)$ and $(0,3)$ MC shells). In the latter case, standard Mackay icosahedra are obtained. According to the estimates of the optimal mismatch values, for Na@K, Na@Rb and Na@K@Rb we expect that the clusters with Ch1 shells are more stable than the Mackay clusters. This is confirmed by the data reported in Supplementary Tables 2-4, where it is shown that Mackay clusters are energetically unstable upon atomic pair exchanges, while the clusters with Ch1 shells are stable. We note that DFT and Gupta results are in good agreement, which supports the validity of the Gupta model for these clusters.
EXCH 3 SHELLS 4 SHELLS
MC@BG MC@MC MC@(BG-Ch1) MC@(MC-MC)
DFT GUPTA DFT GUPTA DFT GUPTA DFT GUPTA
1 +0.240 +0.322 $-0.345$ $-0.086$ +0.164 +0.251 $-0.046$ +0.013
2 +0.223 +0.262 $-0.323$ $-0.079$ +0.253 +0.273 $-0.064$ +0.006
Table 2: DFT and Gupta data for NaK clusters.
In the initial configurations, these clusters always contain a core with MC shells $(0,0)$ and $(0,1)$ made of Na atoms, while the other shells are made of K atoms. For MC@BG, the $(1,1)$ BG shell is added to the core, whereas for MC@BG-Ch1 a further $(1,2)$ shell is added (see Supplementary Figure 13a). For MC@MC, the $(0,2)$ shell is added to the core, whereas for MC@MC-MC a further $(0,3)$ shell is added. Exchanges of type 1 and 2 (see Supplementary Figure 13b,c) are made on the initial configuration. The energy differences (in eV) between the configurations after and before the exchange of the atomic pair are reported. The exchange is favourable for negative differences, unfavourable otherwise.
EXCH 3 SHELLS 4 SHELLS
MC@BG MC@MC MC@(BG-Ch1) MC@(MC-MC)
DFT GUPTA DFT GUPTA DFT GUPTA DFT GUPTA
1 +0.283 +0.381 $-0.711$ $-0.415$ +0.205 +0.284 $-0.081$ $-0.024$
2 +0.237 +0.289 $-0.691$ $-0.395$ +0.293 +0.289 $-0.137$ $-0.045$
Table 3: DFT and Gupta data for NaRb clusters.
The same as in Table 2 but with Rb atoms instead of K atoms. Energies are in eV.
EXCH MC@BG@Ch1 MC@MC@MC
DFT GUPTA DFT GUPTA
1 +0.180 +0.223 $-0.025$ +0.031
2 +0.242 +0.237 $-0.042$ $-0.006$
3 +0.064 +0.071 +0.031 +0.047
4 +0.048 +0.054 +0.019 +0.028
Table 4: DFT and Gupta data for NaKRb clusters.
In the initial configurations, these clusters always contain a core with MC shells $(0,0)$ and $(0,1)$ made of Na atoms, while the third and fourth shells are made of K and Rb atoms, respectively. For MC@BG@Ch1, the $(1,1)$ and the $(1,2)$ shells are added to the core (see Supplementary Figure 13a), whereas for MC@MC@MC the $(0,2)$ and $(0,3)$ shells are added. Exchanges of types 1, 2, 3 and 4 (see Supplementary Figure 13b,e) are made on the initial configuration. The energy differences (in eV) between the configurations after and before the exchange of the atomic pair are reported. The exchange is favourable for negative differences, unfavourable otherwise.
Appendix S3 Growth simulations
Here we present additional results on the simulations of the growth after deposition of atoms one-by-one on a preformed seed [44, 24]. These simulations show a few more examples of growth sequences that can be represented by paths in the hexagonal lattice.
In Supplementary Figure 14, three growth sequences for binary and ternary alkali metal clusters are shown. In Supplementary Figure 14a, we report the deposition of K atoms on a Na@K cluster terminated by a BG $(1,1)$ shell of K atoms. This is a case of deposition without mismatch, therefore the growth continues within the Ch1 class as expected.
In Supplementary Figure 14b, we show the result of the deposition of Cs atoms on a Na@Rb seed. The seed is terminated by a Ch1 $(1,3)$ shell of Rb atoms. According to the path rules, either a $(1,4)$ or a $(2,3)$ shell can form upon deposition of further atoms. The optimal size mismatch of these steps can be estimated by Eq. S53, where the coefficient $\xi$ is calculated recursively by applying Eq. S67 twice, since there are three shells made of the same atom type ($(1,1)$, $(1,2)$ and $(1,3)$). The optimal values of the mismatch are 0.0547 and 0.1162 for the steps towards $(1,4)$ and $(2,3)$, respectively. As usual, the class-conserving step is associated with the lower optimal mismatch. The mismatch between Cs and Rb, as evaluated from their bulk lattice constant, is of 0.077. Since Cs atoms are too large for accommodating on the $(1,4)$ shell, we expect the $(2,3)$ shell to form. This prediction finds confirmation in our growth simulations, where we observe the formation of a perfect Ch2 shell. We note that Cs and Rb atoms show some tendency towards intermixing [53], so that there are a few exchanges of Rb and Cs atoms during growth that bring some Rb atoms to the cluster surface.
In Supplementary Figure 14c a more complex growth sequence is shown. The seed is Na@K terminated by a Ch1 $(1,2)$ shell of K atoms, on which Rb atoms are deposited. Two steps are possible, leading to the formation of either the Ch1 $(1,3)$ shell or the BG $(2,2)$ shell. The optimal values of the mismatch are 0.0618 and 0.1129, respectively. The mismatch between K and Rb is 0.0716, which is slightly larger than the optimal mismatch for the growth of a Ch1 shell. Indeed, as in the case of Supplementary Figure 14b, the chirality class is changed and the BG $(2,2)$ shell is formed. This is followed then by a $(2,3)$ Ch2 shell, i.e. the only viable alternative for the growth on top of the BG shell. However this is not the end of the story, because upon further Rb atom deposition there is first a sudden transformation of the subsurface shell from $(2,2)$ to $(1,3)$ and then of the $(2,3)$ shell from $(2,3)$ to $(1,4)$. These transformations take place by sudden collective reshaping processes, of the type frequently observed in metal clusters [24]. The final result is as if the cluster would have grown into the Ch1 class from the beginning. The transformation of the first Rb shell from $(2,2)$ to $(1,3)$ is due to the packing of multiple shells, which affects the optimal mismatch at the Na-Rb interface. Specifically, the addition of further shells slightly increases the optimal mismatch (see Supplementary Note 1.7). The optimal mismatch for $(1,2)$-$(1,3)$ interface is therefore likely to approach and eventually overcome the mismatch between K and Rb, so that the Ch1 shell becomes more favourable than the BG one. This induces the reshaping of the shell. After the transformation of the first Rb shell, we have a $(2,3)$ Rb shell on top of a $(1,3)$ shell made of atoms of the same type. This configuration is not the optimal one, because when the mismatch is zero atoms are better accommodated on shells of the same chirality class. Therefore, the transformation of the second Rb shell towards the Ch1 $(1,4)$ arrangement takes place.
In Supplementary Figure 15 we show a growth sequence for a alkali metal cluster made of four different elements. The seed is Na@K@Rb terminated by a Ch1 $(1,3)$ shell of Rb atoms, on which Cs atoms are deposited. Two steps are possible, leading to the formation of either the Ch1 $(1,4)$ shell or the Ch2 $(2,3)$ shell. The optimal values of the mismatch are 0.0183 and 0.0778, respectively. We note that these values are different from those of Supplementary Figure 14, referring to the same steps on the closed-packed plane, because here in the seed there is a single shell of Rb atoms instead of three. The mismatch between Rb and Cs is 0.0769, which is very close to the optimal mismatch for the formation of the Ch2 $(2,3)$ shell. Such shell is indeed formed in our simulations. However, the deposition of further Cs atoms leads to the formation of the BG $(3,3)$ shell instead of the Ch2 $(2,4)$ shell, which is what one would expect when atoms of the same species are deposited. However, we note that, in this case, the optimal mismatch for both steps is quite small, being 0.0249 for the class-conserving step and 0.0459 for the class-changing one. As such, even though the Ch2 shell is expected to be more favourable, the BG one is not that unfavourable and therefore may form during the growth simulation. Besides, due to the mixing of Cs atoms with the smaller Rb atoms in the $(2,3)$ shell, the effective mismatch is actually larger then zero, which helps in stabilizing the BG shell. Upon further deposition of Cs atoms, the Ch3 $(3,4)$ shell is formed, corresponding to the natural growth step on top of the BG $(3,3)$ shell.
In Supplementary Figure 16 data on the growth of Ag atoms on Ni and Co Mackay seeds of 147 and 309 atoms are shown. The large mismatch (0.161 for Ag on Ni and 0.156 for Ag on Co) is close to optimal for the growth of Ch1 shells on seeds in this size range; this kind of growth is indeed observed in the growth simulations.
In Supplementary Figure 17 the growth of Ag atoms on a Mackay seeds of size 309 is shown. This cluster is terminated by a $(0,4)$ shell. For the growth of the first Ag shell, the mismatch between Ag and Cu (0.129) is closer to the optimal one of an AM1 shell. The AM1 $(1,4)^{*}$ shell is indeed grown, with very few defects. Then the growth on the AM1 shell continues by a class-changing path by forming defective AM shells up to the AM3 $(1,4)^{*}$ shell. Upon further deposition, there is a sudden transformation of all Ag shells to the Ch1 arrangement, up to the $(1,6)$; this configuration is energetically more favourable.
In Supplementary Figure 18 the growth of Ag on imperfect Mackay Ni seeds is shown. These seeds present an incomplete Mackay external shell. In spite of the defective character of the seed, Ch1 shells are formed. These shell are slightly distorted because of the asymmetries in the starting core, but they very much resemble the perfect Ch1 shells. This shows that the growth mode of chiral shell is quite robust against imperfections.
In Supplementary Figure 19 the growth mechanism of AM shells one on top of the other is discussed. This mechanism is dominated by the trapping of atoms in fourfold adsorption sites on the surface, which naturally leads to class changes at every step in the hexagonal plane.
Figure 14: Growth simulations of alkali metal clusters of two and three different elements.
a Deposition of K atoms on a Na@K seed of 45 atoms, in which a Mackay core of 13 atoms is covered by one Ch1 shell of K atoms. The growth temperature is 125 K and the deposition rate is 1 atoms/ns. Deposited K atoms form two further Ch1 shells. b Deposition of Cs atoms on a Na@Rb seed of 249 atoms, in which a Mackay Na inner core is covered by two Ch1 shells of Rb atoms. The growth temperature is 125 K and the deposition rate is 1 atoms/ns. Cs atoms form a Ch2 shell. c Deposition of Rb atoms on a Na@K seed of 117 atoms, made of a Mackay inner core of 13 Na atoms and two Ch1 shells of K atoms. Growth temperature is 125 K and deposition rate is 1 atoms/ns. Deposited Rb atoms first form the $(2,2)$ BG shell, which is completed at 239 atoms, and then the growth continues by the formation of a $(2,3)$ Ch2 shell. Upon deposition of further Rb atoms, the $(2,2)$ shell suddenly transforms into the $(1,3)$ Ch1 shell, still covered by a $(2,3)$ Ch2 shell, which then transforms into the $(1,4)$ Ch1 shell.
Figure 15: Growth simulations of alkali metal clusters of four different elements.
Deposition of Cs atoms on a Na@K@Rb seed of 249 atoms, in which a Mackay core of 13 atoms is covered by two Ch1 shell of K atoms, and by a Ch1 shell of Rb. The growth temperature is 125 K and the deposition rate is 1 atoms/ns. Deposited Cs atoms form Ch2 and Ch3 shells
Figure 16: Growth sequences for Ag atoms on Ni and on Co seeds.
A Deposition of Ag atoms on a Ni309 Mackay icosahedral seed, at $T=450$ K, and deposition rate 0.1 atoms///ns. B Deposition of Ag atoms on a Co147 Mackay icosahedral seed, at $T=350$ K, and deposition rate 0.1 atoms///ns. c Deposition of Ag atoms on a Co309 Mackay icosahedral seed, at $T=450$ K, and deposition rate 0.1 atom///ns. In A-C, we show the growth path (the part of the path corresponding to the red arrows) in the hexagonal lattice and images of the cluster surface at different growth stages.
Figure 17: Growth sequence for Ag atoms on a Cu seed.
In the top row, the growth sequence is shown for the deposition of Ag atoms on a Cu309 Mackay icosahedral seed, at $T=350$ K, and deposition rate 1 atom///ns up to size 1175. The growth produces AM structures changing class at each step in the hexagonal plane. The bottom row show the continuation of the growth up to size 1265. In this part of the simulation, a sudden rearrangement of the Ag shells to Ch1 structures takes place, so that the final structure does not contain AM shells anymore and its structure is represented by the the path in the bottom row. The sudden rearrangement is due to the very unfavourable energy of the AM multi-shell arrangement, which is highly metastable.
Figure 18: Growth of Ag atoms on a Ni defective seeds.
The growth simulations are at $T=450$ K, and deposition rate 0.1 atoms/ns. At variance with Supplementary Figures 16 and 17, here the images represent different shells of the final structures, obtained at the end of the growth, instead of snapshots taken during the growth. In a growth starts from a Ni289 seed, which is terminated by a slightly incomplete $(0,4)$ MC shell. In the final structure, the $(0,4)$ shell is completed by Ag atoms, then covered by Ch1 shells. In b growth starts from a Ni329 seed, which is terminated by a small fragment of a (0,5) MC shell. in the final structure, that shell is completed by Ag atoms, which transform it into a Ch1 $(1,4)$ shell, which is covered by further Ch1 shells. In c the growth starts from a Ni227 seed, which is terminated by a half $(0,4)$ MC shell. In the final structure, that shell is completed by Ag atoms, so that a the resulting shell presents a Ni and an Ag half. On the Ni side, the shell is MC $(0,4)$, while on the Ag side the shell is Ch1 $(1,3)$. This is indicated by the bifurcation in the path. Further shells are of class Ch1.
Figure 19: Growth mechanism on AM shells.
a The four orange atoms identify a fourfold adsorption site on the $(1,3)^{*}$ AM shell. b A deposited adatom (cyan) is preferentially trapped in a site of this type. c A new fourfold adsorption site is created on top for the four orange atom so that d it can trap a further deposited atom. e-g The process is self-replicating, finally leading to the formation of the $(2,3)^{*}$ AM shell. h The atoms of $(1,4)^{*}$ AM shell on top of a $(1,3)^{*}$ AM shell find only threefold adsorption sites, identified by the orange atoms. Sites at the facet edge are twofold (see the blue atoms) so that they are not even locally stable for isolated adatoms.

### v3

@@ Acknowledgments
Figure 2
Mapping icosahedra into paths. a Path of a Mackay icosahedron: $i=5$ shells are assembled on top of each other. b The three possible cases when choosing between $(h,k)\to(h,k+1)$ and $(h,k)\to(h+1,k)$ at each step of the path. In b, points are colored according to the chirality class of the corresponding shell, as in Figure 1b.
Figure 3
Optimal size mismatch. a Binding energy per atom for clusters made of a core with five MC shells plus a sixth shell, which is either MC (blue curve) or Ch1 (red curve). Interactions are of Lennard-Jones (LJ) type (see Methods). The particles of the sixth shell differ from those in the core only by their size. The energy is in units of the $\varepsilon$ of the LJ potential. The structures are shown in b. c Comparison of the optimal mismatch between steps $(0,k)\to(0,k+1)$ (blue curve) and $(0,k)\to(1,k)$ (red curve), corresponding to additional MC and Ch1 shells on an MC core, respectively. The optimal mismatch of Eq. (3) (GEO values) is compared to LJ and Morse potential results. The Morse potential data are given for three values (4,5,6) of the parameter $\alpha$, that regulates the width of the potential well (see the Methods section). d Comparison between steps $(1,k)\to(1,k+1)$ (yellow curve) and $(1,k)\to(2,k)$ (orange curve), corresponding to additional Ch1 and Ch2 shells on a core containing $k$ MC and one Ch1 shells, respectively. The optimal mismatch of Eq. (3) (GEO) is compared to LJ data. Source data of a, c, d are provided as a Source Data file.
Figure 4
A path with spontaneous symmetry breaking. a A path connecting all BG shells through their nearby chiral shells. Black and grey arrows indicate paths with right and left-handed chiral shells. A generic path of this type may alternate both chiralities, corresponding to combinations of black and grey arrows. Shells are enumerated along the black-arrow path. b Optimal mismatch $\text{sm}_{i,i+1}$ according to Eq. (3) for the path in a. Points connected by blue and red lines correspond to changes of species every shell and every two shells. The optimal values of $\text{sm}_{i,i+1}$ are compared to the mismatch between pairs of atomic species, indicated by squares and diamonds. The triangles correspond to the optimal mismatch between two MC shells and two Ch1 shells. c Structures for $i=2,3,4$ along the path of a, with their compositions related to specific systems. Source data of b are provided as a Source Data file.
Figure 5
DFT data for atomic pair exchanges. Perfect structures before the atomic pair exchange, either of a MC@BG/MC@BG@Ch1 or d MC@MC/MC@MC@MC type. Atoms in the second and third shell are colored in brown and yellow, respectively. b, c, e, f configurations after the atomic pair exchanges of two different types. Energy differences with respect to the perfect configurations in a and d are reported below. Positive and negative energy differences indicate unfavourable and favourable exchange processes, respectively. Complete data are reported in Supplementary Tables 1-3.
Figure 6
Growth sequences of chiral icosahedra. a Rb atoms are deposited on an icosahedral Na@${}_{13}@$Rb32 seed. b The seed is Na@${}_{13}@$Rb32@Rb72, on which Cs atoms are deposited. c Size mismatch for some transition metal pairs compared to the optimal mismatch for a MC and a Ch1 shell on a MC core (brown and green lines, respectively), as a function of core size. d-e Snapshots from growth simulations and corresponding paths for the deposition of Ag atoms on d Ni147 and g Cu561 cores. All simulation snapshots in a-b, d-e are taken at magic sizes for the corresponding paths. In the top and bottom rows of the snapshot sequences we show the cluster surface and its cross section, respectively. In the representation of shell sequences in the hexagonal plane, black arrows are used to connect the shells belonging to the initial icosahedral seed, while red arrows connect the shells spontaneously formed on top of it in the growth simulations. Na, Rb, Cs, Ni, Cu and Ag atoms are colored in brown, yellow, dark green, light green, orange and grey, respectively. Source data of c are provided as a Source Data file.
Figure 7
@@
Figure 8
Growth of generalized anti-Mackay structures. a Optimal mismatch for one AM1 shell on a Mackay core ($(0,k)\to(1,k)^{*}$) as a function of core size (yellow curve), compared with Ch1 ($(0,k)\to(1,k)$) (red curve) and Mackay shells ($(0,k)\to(0,k+1)$) (blue curve), for Lennard-Jones clusters. GEO values are calculated by the formula derived in Supplementary Note 1.5. The mismatch between Ag and Cu is indicated. b,c Snapshots from MD growth simulations (with cluster surfaces and cross sections shown in top and bottom lines, respectively) and corresponding paths, with AM shells indicated by squares. In b Ag atoms are deposited on a Mackay Cu147 core. In c Pd atoms are deposited on a Ni core and then Ag atoms are deposited on the Ni@Pd cluster. In the top and bottom rows of the snapshot sequences we show the cluster surface and its cross section, respectively. In the representation of shell sequences in the hexagonal plane, black arrows are used to connect the shells belonging to the initial icosahedral seed, while red arrows connect the shells spontaneously formed on top of it in the growth simulations. Ni, Cu, Pd and Ag atoms are colored in light green, orange, blue and grey, respectively. Source data of a are provided as a Source Data file.